How many pounds of force would it take for a class three lever to develop 230


pounds of resultant force if applied force arm is 18 and ¼ inches, and the resultant


force arm is 36 and 7/8 inches?

1
Expert's answer
2021-11-11T12:00:05-0500

The force is


"F=f\frac{L}{l}=230\frac{36+7/8}{18+1/4}=465\text{ lbf}"